OpenStudy (scienceteen8):

is 38/53 in simplest form? Explain why or why not. If not, write it in simplest form

OpenStudy (hennawho):

is there a gcf in the numerator and denominator?

OpenStudy (scienceteen8):

no

OpenStudy (aihberkhan):

Okay. go for prime factors top and bottom 38 = 2 * 19 53 = 53 * 1 nothing to cancel in 2 * 19/53 so 38/53 IS as simple as you can write it.

OpenStudy (hennawho):

then its in its simplest form

OpenStudy (aihberkhan):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! @scienceteen8

OpenStudy (scienceteen8):

thnxs